My question asks, “Part of a parabola and part of a line are shown on the coordinate plane below. Will they ever intersect? Support your solve solution with mathematics.”
I know that they will intersect, but how do I support it with mathematics? The solutions that I get involve imaginary numbers, so I need help.
The equations are the following:
y=10x-3
y=(x^2)+15x+55.5
Thank you!

Answer by Alan3354(69443)

y=10x-3
y=(x^2)+15x+55.5
===============
10x-3=(x^2)+15x+55.5
x^2 + 5x + 58.5 = 0
------------
How do you know they will intersect?